# Ultrastable 2D glasses and packings explained by local centrosymmetry

### v1

@@ Abstract
For isostatic packings, it has been shown rigorously [23] that this negative nonaffine contribution exactly cancels the affine contribution, thus leading to the vanishing of rigidity or unjamming at $z_{c}=2d$ for central-force packings at the random close packing (RCP) [26], where $z$ is the coordination number and $d$ the space dimension. Therefore, the statistical degree of centrosymmetry in local force-transmission, introduced in Ref. [11], can be leveraged as a useful metric to predict the mechanical and vibrational properties of a amorphous systems solely based on the static particle-level structure of the system, and this usefulness has been benchmarked several times in the past on experimental data for very different systems such as colloidal glasses [9, 10], jammed packings [20] and metallic glasses [4]. Furthermore, it was shown already in [11] that the inversion-symmetry order parameter $F_{IS}$ not only provides a strong correlation with the mechanical response but also with the vibrational properties, such as the boson peak excess of soft vibrational modes over the Debye prediction. At the same time, a traditional metric such as the bond-orientational order parameter [15] is unable to correlate with the mechanical and vibrational properties of disordered systems such as random packings and random networks [11].
@@
$|\boldsymbol{\Xi}|^{2}=\kappa^{2}R_{0}^{2}\sum_{i}\sum_{\alpha\in\{x,y\}}\left(\sum_{j\in\mathrm{nn}(i)}n_{ij}^{\alpha}\,n_{ij}^{x}\,n_{ij}^{y}\right)^{2}.$ (3)

### v2

@@ Abstract
For isostatic packings, it has been shown rigorously [23] that this negative nonaffine contribution exactly cancels the affine contribution, thus leading to the vanishing of rigidity or unjamming at $z_{c}=2d$ for central-force packings at the random close packing (RCP) [26, 28], where $z$ is the coordination number and $d$ the space dimension. Therefore, the statistical degree of centrosymmetry in local force-transmission, introduced in Ref. [11], can be leveraged as a useful metric to predict the mechanical and vibrational properties of a amorphous systems solely based on the static particle-level structure of the system, and this usefulness has been benchmarked several times in the past on experimental data for very different systems such as colloidal glasses [9, 10], jammed packings [20] and metallic glasses [4]. Furthermore, it was shown already in [11] that the inversion-symmetry order parameter $F_{IS}$ not only provides a strong correlation with the mechanical response but also with the vibrational properties, such as the boson peak excess of soft vibrational modes over the Debye prediction. At the same time, a traditional metric such as the bond-orientational order parameter [15] is unable to correlate with the mechanical and vibrational properties of disordered systems such as random packings and random networks [11].
@@
$|\boldsymbol{\Xi}|^{2}=\kappa^{2}\sum_{i}\sum_{\alpha\in\{x,y\}}\left(\sum_{j\in\mathrm{nn}(i)}R_{ij}\,n_{ij}^{\alpha}\,n_{ij}^{x}\,n_{ij}^{y}\right)^{2}.$ (3)